If the mean GPA among students is 3.25 with a standard deviation of 0.75, and we select a random sample of 300 people, at what value for the sample mean would be greater than exactly 95% of possible sample means?
Answer #1


Solution :

Given that,

\bar x = 3.25

s =0.75

n = 100

Degrees of freedom = df = n - 1 = 100 - 1 = 99

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,99 =1.984

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

=1.984 * (0.75 / \sqrt 100)

= 0.158

The 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

3.25 - 0.158 < \mu < 3.25 + 0.158

3.092 < \mu < 3.409

(3.092, 3.409 )